OpenStudy (anonymous):

find the derivative of csc exponent of -1 3x?

OpenStudy (dumbcow):

\[y = \csc ^{-1}(3x)\] \[\csc y = 3x\] \[\sin y = \frac{1}{3x}\] now differentiate both sides \[y' \cos y = -\frac{1}{3x^{2}}\] \[y' = -\frac{\sec y}{3x^{2}}\]